OpenStudy (anonymous):

When a person swallows giardia cysts, stomach acids and pancreatic enzymes cause the cysts to release trophozoites, which divide every 12 hours. assume that the model is given by an xponential function with base 2. a) Suppose that the number of trophosoites at time t=0 is y0. write a function giving the number after t hours

OpenStudy (anonymous):

at time t=0 is \[y _{0}\]

myininaya (myininaya):

too many sciencey words

OpenStudy (anonymous):

haha i know!

myininaya (myininaya):

y=a*2^t y0=a y=y0*2^t

OpenStudy (anonymous):

just wonderin how you got the 2 from e

myininaya (myininaya):

base is 2 not e

OpenStudy (anonymous):

is the 2 u wrote there the base?

OpenStudy (anonymous):

dw i got it lol

myininaya (myininaya):

dw?

myininaya (myininaya):

whats that mean

OpenStudy (anonymous):

dont worry haha

myininaya (myininaya):

oh lol cool

OpenStudy (anonymous):

So there is a part B to this problem which I am having trouble with: A single trophozoite can multiple to a million in just 10 days and to a billion in 15 days. Verify this fact